# Statistical noise enhances quantumness benefits in spin-network quantum reservoir computing

### v1

@@ Abstract
Quantum reservoir computing offers a promising approach to the utilization of complex quantum dynamics in machine learning. Statistical noise inevitably arises in real settings of quantum reservoir computing (QRC) due to the practical necessity of taking a small to moderate number of measurements. We investigate the effect of statistical noise in spin-network QRC on the possible performance benefits conferred by quantumness. As our measures of quantumness, we employ both quantum entanglement, which we quantify by the partial transpose of the density matrix, and coherence, which we quantify as the sum of the absolute values of the off-diagonal elements of the density matrix. We find that reservoirs which enjoy a finite degree of quantum entanglement and coherence are more stable against the adverse effects of statistical noise on performance compared to their unentangled, incoherent counterparts. Our results indicate that the potential benefit reservoir computers may derive from quantumness depends on the number of measurements used for training and testing, and may indeed be enhanced by statistical noise. These findings not only emphasize the importance of incorporating realistic noise models, but also suggest that the search for quantum advantage may be aided rather than impeded by the practical constraints of implementation within existing machines.
I Introduction
@@
As we advance further into the information age, it is imperative that we should scale up our ability to interact with data of unprecedented complexity and volume. It is thus of increasingly critical importance to endow machines with the ability to understand, manipulate, and respond to sequential data inputs. A primary challenge in this arena arises in the form of the so-called von Neumann bottleneck, in which the physical division of processing and memory units constrains processing speed [1]. In stark contrast, biological systems manage dynamic, continuous information processing with vastly superior efficiency and remarkably low energy usage [2, 3].
@@
Reservoir computing (RC) is as a paradigm that promises to empower our machines in the face of these challenges [4, 5, 6]. At the heart of it is the reservoir, a dynamical system of very high dimensionality, which receives incoming streams of data and naturally generates transient internal states endowed with fading memory and nonlinear processing capabilities. Such dynamical complexity lends itself well to machine learning tasks requiring a strong capacity for retaining prior inputs, such as speech recognition, stock market prediction, and autonomous motor control for robots [7]. Early implementations of RC used randomly connected artificial neural networks or spiking neural networks [8], while physical realizations have emerged across various platforms, such as photonics [9, 10, 11, 12, 13], phonons [14, 15], magnons [16, 17, 18], spintronics [19, 20, 21], and neuromorphic nanomaterials [22, 23].
[h]
@@
Figure 1: Schematic representation of spin-network quantum reservoir computing. The reservoir (red box) comprises $N$ qubits (blue spheres) with random and unchanging connections therebetween. At each timestep, the input data is injected into a fraction of the qubits, true qubits (purple box). Following injection, measurements are collected and sent to a classical decoder. The learning process is undertaken only by the classical connections between the reservoir and the decoder.
@@
Recent years have witnessed the rise of quantum reservoir computing (QRC), which seeks to employ within the framework of RC the unique properties of quantum mechanics, such as quantum entanglement and coherence [24, 25]. One model of QRC utilizes networks of qubits [25, 26, 27], and another employs reservoirs comprising quantum-mechanical oscillators [28, 29, 30]. Here, we focus on the qubit-based approach, in which changes to qubit states are driven by sequential inputs that propagate through the reservoir’s quantum dynamics [31, 32]. Time-series data enters the system by injection into the so-called true nodes. The system, comprising both true nodes and hidden nodes, is allowed to evolve under the influence of the input. Finally, an output signal is read from all nodes, to be used in the learning stage. Crucially, training only takes place at the output layer, which simplifies the learning considerably. A schematic of this general QRC mechanism is shown in Fig. 1.
@@
The chief obstacle in the face of quantum computing and quantum machine learning is the decoherence inherent to the operation of the available quantum devices. Significant efforts have been dedicated to correcting or mitigating the resulting errors. However, there have been indications that, in certain scenarios, this dissipative noise may be harnessed to provide a computational benefit in the context of QRC [33, 34, 35]. Another recent observation of a potential slight enhancement of the performance of weakly-interacting spin-network QRC was attributed to dissipation [36].
@@
There has also been a rising interest in understanding the potential advantage conferred by the essential aspects of quantum mechanics — quantum entanglement and superposition — in QRC. Such an advantage stands in contrast with the rather more traditional advantage of the exponential scaling of the state space, which can be present in the absence of quantumness [37]. Considerable research efforts have endeavored to investigate how quantumness might provide benefit to machine learning, in distributed learning over quantum networks [38], spin-network QRC [39], and oscillator-based QRC [40]. For example, Ref. 36, which sought to understand the physical circumstances conducing to an entanglement advantage in spin-network QRC, demonstrated that the presence of the entanglement advantage was, in the presence of dissipation, dependent on the frequency scale at which the input signal varies. This was interpreted as a consequence of the timescale introduced by dissipation, which determines whether quantum memory in the system can survive for long enough in the system for the input to manifest its temporal features, allowing the quantum reservoir computer to remember them.
@@
Another important frontline in QRC is the ability to implement them on real machines. There are a number of candidate platforms which lend themselves well to the task, such as Rydberg atoms[41], superconducting qubits [42], photonics [13], and trapped ions [43]. Making contact with such implementations, however, faces a number of challenges, such as the necessity of contending with the measurement problem of quantum mechanics. Traditional QRC relies on projective measurements which destroy the quantum state, and thus require rewinding the system and input back in time every time a measurement is made. This, in combination with the necessity of repeating the process many times to compute quantum expectation values, leads to unfavorable time complexity and the need for an external memory. A number of approaches have been proposed to address this, such as those utilizing weak measurements [44], feedback protocols [45], an approach that combines the two [46], and artificial memory restriction [47]. However, it is generally the case that a real implementation will be rather severely constrained in the number of measurements that can be practically extracted from the machine.
@@
In this work, we seek to further understand the consequences of the statistical error introduced by the necessity of making a limited number of measurment in spin-network QRC, inspired by Ref. [48]. We are interested in understanding how this error interacts with the advantages that may be conferred by quantumness, which we measure by means of quantum entanglement and coherence. We find that the advantages potentially offered by quantumness are qualitatively dependent on the strength of this statistical noise; while having an overall adverse effect on performance, statistical noise is less detrimental to reservoir computers with higher amounts of quantum entanglement and coherence than reservoirs which are unentangled and incoherent. Our results not only highlight the importance of accounting for statistical noise in QRC simulations, but also demonstrate that the necessity to make fewer measurements may actually lead QRC systems to derive more benefit from quantumness. Indeed, the presence of this statistical noise was found to give rise to a quantumness benefit where there was none before.
@@
The remainder of this paper is organized as follows: in section II we describe our dynamical models and methodology. We present and discuss our results in section III, and finally outline our conclusions in section IV.
II Model and Methodology
II.1 Physical System
@@
Our quantum reservoir is a network of $N=4$ qubits obeying a generalized transverse-field Ising model [49, 50] with the Hamiltonian
@@
where $\hat{\sigma}_{i}^{a}$ ($a=x,y,z$) are the Pauli operators, $h$ is the transverse magnetic field, and $J_{ij}$ are randomly generated network connectivities sampled from a uniform distribution in the interval $[-J_{s}/2,J_{s}/2]$. Ours is an open quantum system experiencing Markovian dynamics and obeying the Lindblad master equation
@@
$\Gamma$ being the dissipation rate in this high-temperature decoherence channel [51], and $\{\hat{L}_{i}\}$ are identifiable as the raising and lowering operators of the system. The unitary system experiences a dynamical phase transition between an ergodic phase and a many-body-localized phase determined by the ratio $h/J_{s}$ [32].
Figure 2: (left) Examples of input sequences at three different frequency scales $f=0.2$ (top), $f=1$ (middle), $f=5$ (bottom). The red vertical dashed line corresponds to the maximum time delay at which the reservoir is successful in remembering past input. (right) The corresponding outputs of the reservoir at an interaction strength of $J_{s}=1$, a transverse field of $h=2$, an injection period of $\Delta=2.5$, and a dissipation strength of $\Gamma=0.01$. Blue corresponds to the input qubit.
II.2 Input, Training, and Statistical Noise
We utilized input signals $s_{k}$ with a frequency scale $f$, defined as the sum of 20 frequencies $\left\{f_{i}\right\}$ chosen with equal linear spacing in the interval $\left[f/5000,f/50\right]$:
@@ II.2 Input, Training, and Statistical Noise
where $t_{k}$ is the time after $k$ time steps, and $\zeta$ is sampled from the uniform distribution in the interval $\left[0,1\right]$. Input signals are normalized so as to lie between 0 and 1. The left panels of Fig. 2 show examples of these inputs, and the right panels show the corresponding outputs of the quantum reservoir at a representative choice of parameters.
The input signals are introduced into the quantum reservoir by means of reinitializing the state of one qubit that we call qubit 1, a commonly-used form of input injection [52]: we trace out the qubit in question and prepare it in the input-defined state $|\psi_{s_{k}}\rangle=\sqrt{1-s_{k}}|0\rangle+\sqrt{s_{k}}|1\rangle$, thereby transforming the quantum state of the system $\rho$ into the product state
@@
This injection occurs ever $\Delta t=L\delta t$, where $L$ is the number of time steps between successive input injections, chosen so as to lie in the favorable range reported in Ref. [53] in the interest of rich dynamical behavior.
Our reservoirs are time-multiplexed; the outputs are extracted at time intervals of $\Delta t/V$, giving rise to $NV$ virtual nodes. The readout signals are here restricted to $\langle\sigma_{z}\rangle$, and are trained in a linear regression to produce the desired function of the input, which in this work is the fundamental linear memory task [54] $\bar{y}_{k}=s_{k-\tau}$ ($\bar{y}_{k}$ being the target).
@@
The output signals of the reservoir, being a product of our simulations rather than real machines, are subject to a level of Gaussian statistical noise $\sigma$ to simulate the necessity of taking a finite number of measurements when using a real machine as a quantum reservoir compute. The value of $\sigma$, whose effect on the physics is a primary subject of our investigation, corresponds to the number of measurements that would be made in a real setting. As in Ref. [48], the Gaussian standard deviation is taken to be inversely proportional to the square root of the number of measurements.
The quantum reservoir computer is characterized by the fading memory property [55], which guarantees that the memory capacity vanish at large $\tau$. Thus may the total memory capacity of the reservoir computer be defined as
@@
II.3 Entanglement and Coherence
Figure 3: (top) An example of the logarithmic negativity as a function of time - the entanglement abruptly drops whenever the input is injected. Each color corresponds to the bipartition in which the respective qubit labeled in the legend is isolated. (middle) Entanglement vs. interaction strength for the reservoir at a transverse field of $h=2$, and an injection period of $\Delta t=2.5$. The input of frequency of the signal is fixed at $f=0.2$ while the dissipation strength $\Gamma$ is varied. (bottom) Coherence vs. interaction strength for the reservoir at the same physical conditions, at various dissipation strengths and constant frequency.
We measure the amount of quantum entanglement in our reservoir computers by means of the logarithmic negativity [56, 57]
@@ II.3 Entanglement and Coherence
where $\|\cdot\|_{1}$ is the trace norm, $\rho^{\Gamma_{A}}$ is the partial transpose with respect to subsystem A, and all possibe bipartitions of the system are averaged over. Because of our mechanism of input injection, quantum entanglement abruptly plummets each time the input qubit (qubit 1) is reinitialized to reflect the input signal. This drop is naturally strongest for the entanglement corresponding to the bipartition isolating qubit 1. We show an example of this behaviour in Fig. 3 (top panel), where 4 different bipartitions are shown, each isolating one of the 4 qubits.
We quantify the coherence of a quantum state $\rho$ by means of the $l_{1}$-norm of coherence [58, 59], which is given by
@@
and is normalized by its maximum value of $12^{N}-$ 1.
III Results and Discussion
We start by reviewing the behavior of the logarithmic negativity measure of entanglement. In Fig. 3 (middle panel), we present it as a function the interaction strength for a variety of dissipation strengths and input of frequencies, at a transverse field of $h=2$, and an injection period of $\Delta t=2.5$. The system undergoes a dynamical phase transition between an ergodic and many-body localized phase [32], marked by the region between the black, dashed lines. Entanglement peaks to the left of the transition, i.e., in the ergodic regime, and goes to zero at both extremes. It is also worth remarking that the peak is suppressed as $\Gamma$ rises, but the frequency of the input signal has little to no effect.
To contrast, we examine the behavior of the coherence of our reservoirs in Fig. 3 (bottom panel). These results suggest that coherence exhibits a peak as a function of interaction strength only in the case of the isolated resevoir, whereas open systems gain coherence in the many-body localized phase. This behavior is consistent with Lindbladian many-body localization [60]. On the other hand, coherence appears to be insensitive to the frequency of the input, as is entanglement.

### v2

# Youssef Kora 1,2,3,4∗ and Christoph Simon 1,2,3

@@ Abstract
Quantum reservoir computing offers a promising approach to the utilization of complex quantum dynamics in machine learning. Statistical noise inevitably arises in real settings of quantum reservoir computing (QRC) due to the practical necessity of taking a small to moderate number of measurements. We investigate the effect of statistical noise in spin-network QRC on the possible performance benefits conferred by quantumness. As our measures of quantumness, we employ both quantum entanglement, which we quantify by the partial transpose of the density matrix, and coherence, which we quantify as the sum of the absolute values of the off-diagonal elements of the density matrix. We find that reservoirs which enjoy a finite degree of quantum entanglement and coherence are more stable against the adverse effects of statistical noise on performance compared to their unentangled, incoherent counterparts. Our results thus indicate that the potential benefit reservoir computers may derive from quantumness depends on the number of measurements used for training and testing, and that statistical noise, albeit detrimental on the whole, may leave quantum reservoirs in a stronger position relative to less quantum ones. These findings not only emphasize the importance of incorporating realistic noise models, but also suggest that the search for computational regimes that benefit from quantumness may be aided rather than impeded by the practical constraints of implementation within existing machines.
1
[…]
2 Keywords:
quantum reservoir computing, quantum AI, quantum entanglement, statistical noise, spin networks
3 Introduction
@@
As we advance further into the information age, it is imperative that we should scale up our ability to interact with data of unprecedented complexity and volume. It is thus of increasingly critical importance to endow machines with the ability to understand, manipulate, and respond to sequential data inputs with exceeding efficiency. A primary challenge in this arena arises in the form of the so-called von Neumann bottleneck, in which the physical division of processing and memory units constrains processing speed von_Neumann_1993. In stark contrast, biological systems manage, by integrating processing and memory units, dynamic, continuous information processing with vastly superior efficiency and remarkably low energy usage Eliasmith_2012; Stewart_2012.
@@
Reservoir computing (RC) is as a paradigm that promises to empower our machines in the face of these challenges jaeger_2004; maass_2002; verstraeten_2007. At the heart of it is the reservoir, a dynamical system of very high dimensionality, which receives incoming streams of data and naturally generates transient internal states endowed with fading memory and non-linear processing capabilities. Such dynamical complexity lends itself well to machine learning tasks requiring a strong capacity for retaining prior inputs, such as speech recognition, stock market prediction, and autonomous motor control for robots Tanaka_2019. Early implementations of RC used randomly connected artificial neural networks or spiking neural networks Nicola_2017, while physical realizations have emerged across various platforms, such as photonics Vandoorne_2014; Antonik_2017; Larger_2017; Sunada_2021; Garc_2023, phonons Dion_2018; Meffan_2023, magnons Papp_2021; Gartside_2022; K_rber_2023, spintronics Torrejon_2017; Furuta_2018; Tsunegi_2018, and neuromorphic nanomaterials Stieg_2011; Yaremkevich_2023.
@@
Figure 1: Schematic representation of spin-network quantum reservoir computing. The reservoir comprises NN qubits with random and unchanging connections therebetween. At each timestep, the input data is injected into a fraction of the qubits, true qubits. Following injection, a finite number of measurements is collected, which gives rise to statistical noise. The next stage is the learning process, which is undertaken only by the classical connections between the reservoir and the decoder.
@@
Recent years have witnessed the rise of quantum reservoir computing (QRC), which seeks to employ within the framework of RC the unique properties of quantum mechanics, such as quantum entanglement and coherence fujii_2017; fujii_2021. One model of QRC utilizes networks of qubits fujii_2021; Luchnikov_2019; martinez_2020, and another employs reservoirs comprising quantum-mechanical oscillators Govia_2021; Nokkala_2021; Dudas_2023. Here, we focus on the qubit-based approach, in which changes to qubit states are driven by sequential inputs that propagate through the reservoir’s quantum dynamics Nakajima2019; martinez_2021. Time-series data enters the system by injection into the so-called true nodes. The system, comprising both true nodes and hidden nodes, is allowed to evolve under the influence of the input. Finally, an output signal is read from all nodes, to be used in the learning stage. Crucially, training only takes place at the output layer, which simplifies the learning considerably. A schematic of this general QRC mechanism is shown in Fig. 1.
@@
The chief obstacle in the face of quantum computing and quantum machine learning is the decoherence inherent to the operation of the available quantum devices. Significant efforts have been dedicated to correcting or mitigating the resulting errors. However, there have been indications that, in certain scenarios, this dissipative noise may be harnessed to provide a computational benefit in the context of QRC Fry2023; suzuki2022; Domingo2023; sannia_2024. Another recent observation of a potential slight enhancement of the performance of weakly-interacting spin-network QRC was attributed to dissipation kora_2024.
@@
There has also been a rising interest in understanding the potential advantage conferred by the essential aspects of quantum mechanics — quantum entanglement and superposition — in QRC. Such an advantage stands in contrast with the rather more traditional advantage of the exponential scaling of the state space, which can be present in the absence of quantumness ehlers_2025. Considerable research efforts have endeavored to investigate how quantumness might provide benefit to machine learning, in distributed learning over quantum networks gilboa_2023, spin-network QRC gotting_2023, and oscillator-based QRC motamedi_2023. For example, Ref. kora_2024, which sought to understand the physical circumstances conducing to an entanglement advantage in spin-network QRC, demonstrated that the presence of the entanglement advantage was, in the presence of dissipation, dependent on the frequency scale at which the input signal varies. This was interpreted as a consequence of the timescale introduced by dissipation, which determines whether quantum memory in the system can survive for long enough in the system for the input to manifest its temporal features, allowing the quantum reservoir computer to remember them.
@@
Another important frontline in QRC is the ability to implement them on real machines. There are a number of candidate platforms which lend themselves well to the task, such as Rydberg atoms araiza_2022, superconducting qubits suzuki_2022, photonics Garc_2023, and trapped ions haffner_2008. Making contact with such implementations, however, faces a number of challenges, such as the necessity of contending with the measurement problem of quantum mechanics. Traditional QRC relies on projective measurements which destroy the quantum state, and thus require rewinding the system and input back in time every time a measurement is made. This, in combination with the necessity of repeating the process many times to compute quantum expectation values, leads to unfavorable time complexity and the need for an external memory. A number of approaches have been proposed to address this, such as those utilizing weak measurements Mujal_2023; franceschetto_2024, feedback protocols kobayashi_2024, an approach that combines the two monomi_2025, and artificial memory restriction cindrak_2024. However, it is generally the case that a real implementation will be rather severely constrained in the number of measurements that can be practically extracted from the machine.
@@
In this work, we set out to explore the consequences of the statistical error introduced by the necessity of making a limited number of measurement in spin-network QRC, inspired by Ref. palacios_2024. Given certain conditions in which quantumness can confer computational benefits kora_2024, we are now interested in understanding how this error interacts with these benefits. We find that these potential benefits offered by quantumness (which we measure through entanglement and coherence) are qualitatively dependent on the strength of this statistical noise; while having an overall adverse effect on performance, statistical noise is less detrimental to reservoir computers with higher amounts of quantum entanglement and coherence than reservoirs which are unentangled and incoherent. Our results not only highlight the importance of accounting for statistical noise in QRC simulations, but also demonstrate that the necessity of collecting fewer measurements may actually lead QRC systems to derive a relative benefit from quantumness. Indeed, the presence of this statistical noise was found, under certain conditions, to turn a negative relationship between performance and quantumness into a positive one.
The effect of noise in the context of quantum machine learning has been explored in prior studies, such as Ref. fangjun_2023. In that work, they establish general upper bounds on the resolvable expressive capacity under sampling noise. Our study complements this by investigating a concrete spin‐network quantum reservoir (transverse‐field Ising), evaluating memory‐capacity performance, and demonstrating that increasing quantumness (entanglement and coherence) enhances resilience to measurement noise; we observe that under moderate levels of statistical noise, a negative trend between performance and quantumness can transform into a positive one; a ‘noise‐enabled’ effect. Thus does our work offer a distinct and complementary contribution, by explicitly linking quantum resource metrics with noise‐robust reservoir computing performance.
@@
The remainder of this paper is organized as follows: in section 4 we describe our dynamical models and methodology. We present and discuss our results in section 5, and finally outline our conclusions in section 6.
4 Model and Methodology
4.1 Physical System
@@
Our physical model of the quantum reservoir follows our previous formulation in Ref. kora_2024. Figures 1 and 2 are reproduced from that work for clarity. It is a network of $N=4$ qubits obeying a generalized transverse-field Ising model stinchcombe_1973; pfeuty_1971 with the Hamiltonian
@@
where $\hat{\sigma}_{i}^{a}$ ($a=x,y,z$) are the Pauli operators, $h$ is the transverse magnetic field, and $J_{ij}$ are randomly generated network connectivities sampled from a uniform distribution in the interval $J_{s}/2,J_{s}/2]$. Ours is an open quantum system experiencing Markovian dynamics and obeying the Lindblad master equation
@@
$\Gamma$ being the dissipation rate in this high-temperature decoherence channel breuer_2002, and $\{\hat{L}_{i}\}$ are identifiable as the raising and lowering operators of the system. The unitary system experiences a dynamical phase transition between an ergodic phase and a many-body-localized phase determined by the ratio $h/J_{s}$ martinez_2021.
Figure 2: (left) Examples of input sequences at three different frequency scales $f=0.2$ (top), $f=1$ (middle), $f=5$ (bottom). The red vertical dashed line corresponds to the maximum time delay at which the reservoir is successful in remembering past input. (right) The corresponding outputs of the reservoir at an interaction strength of $J_{s}=1$, a transverse field of $h=2$, an injection period of $\Delta=2.5$, and a dissipation strength of $\Gamma=0.01$. Blue corresponds to the input qubit. Reproduced from Ref. kora_2024(© 2024 American Physical Society), with permission.
4.2 Input, Training, and Statistical Noise
As in Ref. kora_2024, we utilized input signals $s_{k}$ with a frequency scale $f$, defined as the sum of 20 signals of equal amplitude and frequencies $\left\{f_{i}\right\}$ chosen with equal linear spacing in the interval $\left[f/5000,f/50\right]$:
@@ 4.2 Input, Training, and Statistical Noise
where $t_{k}$ is the time after $k$ time steps, and $\zeta$ is sampled from the uniform distribution in the interval $\left[0,1\right]$. Input signals are normalized so as to lie between 0 and 1. The left panels of Fig. 2 show examples of these inputs, and the right panels show the corresponding outputs of the quantum reservoir at a representative choice of parameters. This form of the input signal allows us to tune its frequency scale, which is conducive to our goal of explicitly addressing the interplay between the input frequency and the dissipation time scale discussed in Ref. kora_2024.
The input signals are introduced into the quantum reservoir by means of reinitializing the state of one qubit that we call qubit 1, a commonly-used form of input injection mujal_2021: we trace out the qubit in question and prepare it in the input-defined state $|\psi_{s_{k}}\rangle=\sqrt{1-s_{k}}|0\rangle+\sqrt{s_{k}}|1\rangle$, thereby transforming the quantum state of the system $\rho$ into the product state
@@
This injection occurs ever $\Delta t=L\delta t$, where $L$ is the number of time steps between successive input injections, chosen so as to lie in the favorable range reported in Ref. martinez_2020 in the interest of rich dynamical behavior (here taken to be $\Delta t$ = 2.5 and $L$ = 100).
Our reservoirs are time-multiplexed; the outputs are extracted at time intervals of $\Delta t/V$, giving rise to $NV$ virtual nodes (here $V=10$). The readout signals are here restricted to $\langle\sigma_{z}\rangle$, and are trained in a linear regression to produce the desired function of the input, which in this work is the fundamental linear memory task carroll_2022 $\bar{y}_{k}=s_{k-\tau}$ ($\bar{y}_{k}$ being the target).
@@
The output signals of the reservoir, being a product of our simulations rather than real machines, are subjected to a level of Gaussian statistical noise $\sigma$ to simulate the necessity of taking a finite number of measurements when using a real machine as a quantum reservoir computer. The value of $\sigma$, whose effect on the physics is a primary subject of our investigation, corresponds to the number of measurements that would be made in a real setting. As in Ref. palacios_2024, the Gaussian standard deviation is taken to be inversely proportional to the square root of the number of measurements , i.e., $\propto 1/\sqrt{N}$. Thus, for example, $\sigma=0.001$ would represent the physical situation of taking a million measurements.
The quantum reservoir computer is characterized by the fading memory property dambre_2012, which guarantees that the memory capacity vanish at large $\tau$. Thus, the total memory capacity of the reservoir computer may be defined as
@@
Memory capacities are typically computed for inputs which are independent and identically drawn from a probability distribution (i.i.d.), in which case the total memory capacity is bounded by the number of linearly independent state variables of the system dambre_2012. Our input sequences are not i.i.d., but contain periodic structures (as do most real-world sequences), and are thus not subject to that upper bound. On that account, a numerical comparison between the total memory capacity of such periodic sequences and standard i.i.d. benchmarks must be carried out with care. For instance, in Ref. kora_2024, comparisons between capacities of different signals are undertaken by normalizing by the frequency scale.
4.3 Entanglement and Coherence
Figure 3: (top) Entanglement vs. interaction strength for the reservoir at a transverse field of $h=2$, and an injection period of $\Delta t=2.5$. The input of frequency of the signal is fixed at $f=0.2$ while the dissipation strength $\Gamma$ is varied. (bottom) Coherence vs. interaction strength for the reservoir at the same physical conditions, at various dissipation strengths and constant frequency.
We produce reservoir computers of varying entanglement and coherence by varying the interaction strength $J_{s}$, and measure the amount of quantum entanglement therein by means of the logarithmic negativity vidal_2002; plenio_2005
@@ 4.3 Entanglement and Coherence
where $\|\cdot\|_{1}$ is the trace norm, $\rho^{\Gamma_{A}}$ is the partial transpose with respect to subsystem A, and all possibe bipartitions of the system are averaged over.
We have chosen to employ logarithmic negativity as our chief entanglement measure because it is a computable entanglement monotone that is well suited to mixed states and open quantum systems, requiring only the partial transpose of the density matrix. It is thus a widely used and numerically tractable indicator of bipartite quantum correlations in dynamical systems such as the one considered here. For an outlook towards more nuanced alternatives, see Sec. 6.
Because of our mechanism of input injection, quantum entanglement abruptly plummets each time the input qubit (qubit 1) is reinitialized to reflect the input signal. This drop is naturally strongest for the entanglement corresponding to the bipartition isolating qubit 1. We show an example of this behaviour in Fig. 3 (top panel), where 4 different bipartitions are shown, each isolating one of the 4 qubits.
As our secondary measure of quantumness, we quantify the coherence of a quantum state $\rho$ by means of the $l_{1}$-norm of coherence streltsov_2017; baumgratz_2014, which is given by
@@
and is normalized by its maximum value of $2^{N}-1$.
5 Results and Discussion
We start by reviewing the behavior of the logarithmic negativity measure of entanglement. In Fig. 3 (middle panel), we present it as a function the interaction strength for a variety of dissipation strengths, at a transverse field of $h=2$, and an injection period of $\Delta t=2.5$. The system undergoes a dynamical phase transition between an ergodic and many-body localized phase martinez_2021, marked by the region between the black, dashed lines. Entanglement peaks to the left of the transition, i.e., in the ergodic regime, and goes to zero at both extremes. It is also worth remarking that the peak is suppressed as $\Gamma$ rises, but the frequency of the input signal has little to no effect.
To contrast, we examine the behavior of the coherence of our reservoirs in Fig. 3 (bottom panel). These results suggest that coherence exhibits a peak as a function of interaction strength only in the case of the isolated reservoir, whereas open systems gain coherence in the many-body localized phase. This behavior is consistent with Lindbladian many-body localization hamazaki_2022, in which interactions may prevent many-body decoherence. In such a regime, long-range entanglement is degraded by the environment, while coherence remains strong. On the other hand, coherence appears to be insensitive to the frequency of the input, as is entanglement.
Fig. 3 illustrates that quantumness measures such as entanglement and coherence are determined by our choice of the parameters that govern the system’s dynamics. The presence of such quantumness measures can correspond to better performance, depending, as shown in Ref. kora_2024, on the interplay between the dissipation time scale and the input frequency scale. In Ref. kora_2024, such a phenomenon was referred to as an entanglement advantage. We henceforth use that term (entanglement advantage) to refer to a situation where entangled reservoirs exhibit better performance at the task at hand, compared to the same reservoir with low entanglement.